### v2

# Matchings in Matroids over Abelian Groups, III

Abstract.
This paper develops matroidal analogues of classical results on matchings in abelian groups. By embedding matroid ground sets in an abelian group, we introduce base matchings between matroid bases, recover the group-theoretic setting in the uniform matroid case, and derive structural and combinatorial criteria for their existence. Our main focus is on paving matroids. We prove self-matchability for paving matroids, extend asymmetric matchability results using the hyperplane-nullity parameter, and show that stressed hyperplanes provide a natural route to matchability through relaxation.
1. Introduction
1.1. Overview
In an abelian group $G$, a matching is a bijection $f\colon A\to B$ between finite subsets $A,B\subseteq G$ such that $a+f(a)\notin A$ for all $a\in A$. We say that $G$ has the matching property if, for every pair of finite subsets $A,B\subseteq G$ with $|A|=|B|$ and $0\notin B$, there exists such a matching. A geometric model for a class of bipartite graphs was introduced in [11]. In that work, a special kind of perfect matching, called an acyclic matching, was defined and shown to exist for a subclass of these graphs by geometric methods. The existence of acyclic matchings implies the nonvanishing of the determinant of certain weighted biadjacency matrices. This matching framework was later applied to a question posed by E. K. Wakeford in 1920 [25], concerning which sets of monomials can be eliminated from a generic homogeneous polynomial by linear changes of variables. In a special case related to canonical forms of symmetric tensors, Fan and Losonczy reduced Wakeford’s problem to the existence of acyclic matchings in $\mathbb{Z}^{n}$. This property was established for $\mathbb{Z}^{n}$ by Alon, Fan, Kleitman, and Losonczy [7], further developed in [1, 4], and later completely characterized for all abelian groups in [3].
Matchings were extended to the abelian group setting in [18] and to non-abelian groups in [9], with further developments, particularly on enumerative aspects, presented in [14]. A linear formulation was later introduced in [10], and a matroidal analogue was proposed in [6]. Additional progress on matchable matroids appears in [5]. In particular, [6, 5] investigated matchability for the following four classes of matroids:
  • sparse paving matroids,
  • transversal matroids,
  • panhandle matroids,
  • Schubert matroids.
In this paper, we continue this line of research by studying the matchability of paving matroids. Following the terminology introduced in [18], we present the definition of matchings in abelian groups.
Matchings in abelian groups. Throughout, we use the notation $n]:=\{1,\dots,n\}$ for any positive integer $n$. We also write $(G,+)$ for an abelian group with neutral element 0. Let $p(G)$ represent the smallest cardinality of a nontrivial (i.e., non-zero) finite subgroup of $G$. If $G$ has no nontrivial finite subgroup, we make the convention $p(G)=\infty$. Let $A$ and $B$ be finite subsets of $G$ with equal cardinality, such that $0\notin B$. A matching from $A$ to $B$ is a bijection $f:A\to B$ satisfying the condition that $a+f(a)\notin A$ for every $a\in A$. The conditions $|A|=|B|$ and $0\notin B$ are necessary for the existence of such a bijection. A matching is called symmetric if $A=B$, and asymmetric otherwise. If there exists a matching from $A$ to $B$, then $A$ is said to be matched or matchable to $B$. An abelian group $G$ is said to satisfy the matching property if for every pair of finite subsets $A,B\subseteq G$ with $|A|=|B|$ and $0\notin B$, the set $A$ is matched to $B$.
Characterizations of abelian groups satisfying the matching property, as well as necessary and sufficient conditions for the existence of symmetric matchings, were obtained by Losonczy.
Theorem 1.1 ([18]).
Let $G$ be an abelian group, and let $A$ be a nonempty finite subset of $G$. Then there exists a matching from $A$ to itself if and only if $0\notin A$.
Theorem 1.2 ([18]).
An abelian group $G$ satisfies the matching property if and only if $G$ is torsion-free or cyclic of prime order.
Sufficient conditions for the existence of matchings between small enough finite sets in abelian groups were provided in [2].
Theorem 1.3 ([2]).
Let $G$ be an abelian group, and let $A$ and $B$ be finite subsets of $G$ with $|A|=|B|<p(G)$ and $0\notin B$. Then there exists a matching from $A$ to $B$.
A connection to matchings in graphs. While there are meaningful connections between the notion of matchings explored in this paper and the classical matching theory in graphs introduced by Philip Hall in 1935, the two concepts are fundamentally distinct. In particular, in order to address a linear algebra problem posed by Wakeford [25], Fan [11] constructed a bipartite graph
$\mathcal{G}=(V(\mathcal{G}),E(\mathcal{G}))$
from two sets $A,B\subseteq\mathbb{Z}^{n}$. To each element $a\in A$ one associates a symbol $x_{a}$, and to each element $b\in B$ a symbol $y_{b}$. Let
$X=\{x_{a}\mid a\in A\}\qquad\text{and}\qquad Y=\{y_{b}\mid b\in B\}.$
The graph $\mathcal{G}$ is then defined to have vertex set
$V(\mathcal{G})=X\cup Y,$
where an edge joins $x_{a}\in X$ to $y_{b}\in Y$ if and only if
$a+b\notin A.$
In graph theory, a matching is a set of edges with no shared endpoints, and a perfect matching is one that covers all vertices. In the bipartite graph with bipartition sets $V(\mathcal{G})=X\cup Y$, indexed by $A$ and $B$, perfect matchings correspond to bijections $f:A\to B$ such that $a+f(a)\notin A$ for every $a\in A$. Wakeford’s problem [25] thus becomes equivalent to determining the existence of such a bijection, referred to as an acyclic matching from $A$ to $B$.
It is worth noting that, in the classical graph-theoretic setting, a perfect matching merely requires that every vertex is incident to exactly one edge, without reference to a bijection $f:A\to B$ or to the additional condition $a+f(a)\notin A$. By contrast, the notion of matchings in abelian groups incorporates this extra structure in order to capture the linear-algebraic framework underlying Wakeford’s problem.
1.2. Matching in Matroids
We begin with the definition of a matroid. The reader is referred to [20] for all undefined terms. A matroid $M$ is a pair $(E,\mathcal{I})$ where $E=E(M)$ is a finite ground set and $\mathcal{I}$ is a family of subsets of $E$, called independent sets, satisfying the following conditions:
  • $\emptyset\in\mathcal{I}$.
  • If $X\in\mathcal{I}$ and $Y\subseteq X$ then $Y\in\mathcal{I}$.
  • The augmentation property: If $X,Y\in\mathcal{I}$ and $|X|>|Y|$ then there exists $x\in X\setminus Y$ so that $Y\cup\{x\}\in\mathcal{I}$.
Let $M=(E,\mathcal{I})$ be a matroid. The rank of a subset $X\subseteq E$ is given by
$r_{M}(X)=r(X)=max\{|X\cap I|:I\in\mathcal{I}\}.$
The rank of a matroid $M$, denoted by $r(M)$, is defined to be $r_{M}(E(M))$. A set $X\subseteq E$ is called dependent if it is not independent. A maximal independent set is called a basis. It follows from the augmentation property that every two bases have the same size. A minimal dependent set is called a circuit. An element that forms a single-element circuit of $M$ is called a loop. A matroid is called loopless if it does not have a loop.
Given a matroid $M=(E,\mathcal{I})$, the dual matroid $M^{*}=(E,\mathcal{I}^{*})$ is defined so that the bases in $M^{*}$ are exactly the complements of the bases in $M$. A matroid $M=(E,\mathcal{I})$ of rank $n$ is said to be a paving matroid if every (n1)(n-1)-subset of $E$ is independent. If $M$ and $M^{*}$ are both paving matroids, then $M$ is called a sparse paving matroid.
A flat in $M$ is a set $F\subseteq E$ with the property that adjoining any new element to $F$ strictly increases its rank. A flat of rank r(M)1r(M)-1 is called a hyperplane. A set $H\subseteq E(M)$ is called a circuit-hyperplane of $M$ if it is at the same time a circuit and a hyperplane. For any $X\subseteq E$ the difference |X|r(X)|X|-r(X) is called the nullity of $X$ and is denoted by ${\it{null}}(X)$. Let us denote the set of all hyperplanes of $M$ by $\mathcal{H}_{M}$. Then the hyperplane nullity of $M$ is denoted by ${\it{null}}(\mathcal{H}_{M})$ and defined as
${\it{null}}(\mathcal{H}_{M})=\max\{{\it{null}}(H):H\in\mathcal{H}_{M}\}.$
For integers $m$ and $n$ with $0\leq n\leq m$, the uniform matroid $U_{n,m}$ is defined on a ground set $E$ of size $m$ such that a subset $A\subseteq E$ is independent if and only if $|A|\leq n$. Equivalently, the bases of $U_{n,m}$ are exactly the subsets of $E$ of size $n$. Note that uniform matroids are a very special class of sparse paving matroids. A uniform matroid $U_{n,m}$ is called a free matroid if $m=n$.
If $X\subseteq E$, the closure of $X$ is denoted by ${\it{cl}}(X)$ and defined as follows:
${\it{cl}}(X)=\{x\in E:r(X)=r(X\cup\{x\})\}.$
The closure operator satisfies the following properties:
  • For all subsets $X$ of $E$, $X\subseteq{\it{cl}}(X)$.
  • If $X$ and $Y$ are subsets of $E$ with $X\subseteq Y$, then ${\it{cl}}(X)\subseteq{\it{cl}}(Y)$.
  • If r(X)=r(M)1r(X)=r(M)-1, then ${\it{cl}}(X)$ is a hyperplane of $M$.
Following [6] we introduce the definition of matchings in matroids over an abelian group. We may assume that all matroids are loopless (see Remark 5.1 in [6].) We say that $M=(E,\mathcal{I})$ is a matroid over $G$ if $E$ is a subset of $G$.
Definition 1.4.
Let $(G,+)$ be an abelian group.
  1. (1)
    Let $M$ and $N$ be two matroids over $G$ with $r(M)=r(N)=n>0$. Let $\mathcal{M}=\{a_{1},\ldots,a_{n}\}$ and $\mathcal{N}=\{b_{1},\ldots,b_{n}\}$ be ordered bases of $M$ and $N$, respectively. We say $\mathcal{M}$ is matched to $\mathcal{N}$ if $a_{i}+b_{i}\notin E(M)$, for all $i\in[n]$.
  2. (2)
    We say that $M$ is matched to $N$ if for every basis $\mathcal{M}$ of $M$ there exists a basis $\mathcal{N}$ of $N$ such that $\mathcal{M}$ is matched to $\mathcal{N}$.
Note that adopting the notation from Definition 1.4, suppose $\mathcal{M}$ is matched to $\mathcal{N}$ as matroid bases. Then for each $1\leq i\leq n$, we have $a_{i}+b_{i}\notin E(M)$, and in particular $a_{i}+b_{i}\notin\mathcal{M}$. This implies that the map $a_{i}\mapsto b_{i}$ defines a matching in the group-theoretic sense between the subsets $\mathcal{M}$ and $\mathcal{N}$ of $G$. In this way, the definition of matchings in matroids is compatible with the classical notion of matchings in abelian groups.
Compatibility of matchings in matroids with the vector space setting. Definition 1.4 is consistent with the notion of matchability in vector spaces over field extensions, as introduced by Eliahou and Lecouvey [10]. Let $K\subseteq F$ be a field extension, and let $A,B\subseteq F$ be $n$-dimensional $K$-vector subspaces. Suppose $\mathcal{A}=\{a_{1},\dots,a_{n}\}$ and $\mathcal{B}=\{b_{1},\dots,b_{n}\}$ are ordered bases of $A$ and $B$, respectively. Then $\mathcal{A}$ is said to be matched to $\mathcal{B}$ if
(1) $a_{i}^{-1}A\cap B\subseteq\langle b_{1},\dots,\hat{b}_{i},\dots,b_{n}\rangle\quad\text{for every }i\in[n],$
where $\langle b_{1},\dots,\hat{b}_{i},\dots,b_{n}\rangle$ denotes the $K$-vector subspace spanned by $\mathcal{B}\setminus\{b_{i}\}$. The space $A$ is said to be matched to $B$ if every basis $\mathcal{A}$ of $A$ is matched to some basis $\mathcal{B}$ of $B$ in this way.
Now let $G=F\setminus\{0\}$ be the multiplicative group of $F$, and define matroids over $G$ by setting $M=(A\setminus\{0\},\mathcal{I})$ and $N=(B\setminus\{0\},\mathcal{I}^{\prime})$, where $\mathcal{I}$ and $\mathcal{I}^{\prime}$ are the families of linearly independent subsets of $A$ and $B$, respectively. Then
$r(M)=\dim_{K}A=n=\dim_{K}B=r(N).$
Suppose the basis $\mathcal{A}$ is matched to $\mathcal{B}$ in the vector space setting. It follows from condition (1) that $a_{i}b_{i}\notin A$ for all $i$, and hence $A$ is also matched to $B$ in the matroid-theoretic sense.
In this paper we broaden the scope of matroid matchability in three directions. First, we extend the symmetric paving case by proving that any paving matroid is matched to itself under suitable size conditions (Theorem 3.2). Second, we develop an asymmetric extension in which matchability is controlled by the hyperplane-nullity parameter $t$, generalizing the sparse paving setting (Theorem 3.10). Finally, we show that stressed hyperplanes provide a natural route to matchability in relaxations of paving matroids (Theorem 3.23).
The point of passing from sparse paving to paving matroids is not merely to enlarge the class of matroids for which matching results are known. In sparse paving matroids (with rank $n$), the only dependent $n$-sets are circuit-hyperplanes, so the obstruction is essentially binary. In other words, the failure of an $n$-subset to be a basis is controlled by a single type of obstruction, namely circuit-hyperplanes. By contrast, in paving matroids, hyperplanes may have larger nullity, and our results show that matchability is controlled by this richer geometric parameter. In this way, the paper shows that the key mechanism governing matroid matchings is the structure of hyperplanes, rather than near-uniformity alone.
Organization of the paper. Section 2 reviews preliminaries on paving and sparse paving matroids, including their characterization via partitions and hyperplane nullity, as well as additive tools such as Kneser’s theorem and results on critical pairs. Section 3 presents the main theorems: Subsection 3.1 treats the symmetric case, Subsection 3.2 develops the asymmetric theory incorporating hyperplane nullity, and Subsection 3.3 introduces stressed hyperplanes and relaxation, connecting paving matroids to uniform matroids and demonstrating how these ideas ensure broader matchability.
2. Preliminaries
2.1. Notions and Results on Paving Matroids
A $d$-partition of a finite set $E$ is a collection $\mathcal{S}$ of subsets of $E$, all of which are of size at least $d$, so that the intersection of any two distinct sets in $\mathcal{S}$ is of size at most d1d-1. The set $\mathcal{S}=\{E\}$ is a $d$-partition for any $d$, which we call a trivial $d$-partition. The connection between paving matroids and d-partitions is given by the following result, whose proof can be found in [20].
Theorem 2.1.
[20] If $M$ is a paving matroid of rank $n\geq 2$, then its hyperplanes form a non-trivial (n1)(n-1)-partition of $E(M)$. That is, the intersection of any two distinct hyperplanes in $M$ is of size at most n2n-2.
The following result provides an alternative characterization of sparse paving matroids.
Theorem 2.2.
[20] A matroid $M$ of rank $n$ is a sparse paving matroid if and only if every $n$-subset of $E(M)$ is either a basis or a circuit-hyperplane.
Returning to the nullity of hyperplanes, note that if $M$ is a uniform matroid, then clearly $\it{null}(\mathcal{H}_{M})=0$. In other words, hyperplane nullity measures how far a matroid deviates from being uniform. In the more general case where $\it{null}(\mathcal{H}_{M})\leq 1$, we obtain the following characterization, which follows directly from Theorem 2.2 and highlights the role of hyperplane nullity in identifying sparse paving matroids.
Lemma 2.3.
A paving matroid $M$ is sparse paving if and only if ${\it{null}(\mathcal{H}_{M})}\leq 1$.
Proof.
Suppose first that $M$ is a sparse paving matroid. Let $H$ be a hyperplane of $M$. Clearly $|H|\geq n-1$. If |H|=n1|H|=n-1, then $H$ is independent and hence ${\it null}(H)=0$. If |H|>n1|H|>n-1, then $H$ contains an $n$-subset $H^{\prime}$. By Theorem 2.2, $H^{\prime}$ is itself a hyperplane, and thus $H=H^{\prime}$, implying $|H|=n$. In this case, ${\it null}(H)=1$. Therefore, in either case we have ${\it null}(\mathcal{H}_{M})\leq 1$.
Conversely, assume that ${\it null}(\mathcal{H}_{M})\leq 1$. Let $A$ be an $n$-subset of $E(M)$ that is not a basis. We claim that $A$ is a circuit-hyperplane. Since $M$ is paving, then $A$ is a circuit with r(A)=n1r(A)=n-1. Consider the hyperplane $H=\operatorname{cl}(A)$. Then ${\it null}(H)\leq 1$. We have
$1=|A|-r(A)\leq|H|-r(H)={\it null}(H)\leq{\it null}(\mathcal{H}_{M})\leq 1.$
Hence, all inequalities are equalities, which forces $A=H$. Thus $A$ is a hyperplane. Since $A$ is a circuit-hyperplane, by Theorem 2.2, $M$ is sparse paving. ∎
Corollary 2.4.
Let $M$ be a paving matroid. Then
  1. (1)
    $M$ is uniform if and only if ${\it{null}}(\mathcal{H}_{M})=0$.
  2. (2)
    $M$ is non-uniform and sparse paving if and only if ${\it{null}}(\mathcal{H}_{M})=1$.
Proof.
It is immediately seen by invoking Lemma 2.3. ∎
2.2. Abelian Additive Theory
The main additive number theory tools used in the proofs of our main results are Theorem 2.5 and Lemma 2.8.
Let $G$ be an abelian group and $A,B\subseteq G$. The sumset $A+B$ is
$A+B=\{\,a+b:a\in A,\;b\in B\,\}.$
We start with a result of Kneser[19, Theorem 4.3] providing a lower bound on the size of $A+B$.
Theorem 2.5.
For an abelian group $G$ and finite nonempty subsets $A,B\subseteq G$, the sumset $A+B$ satisfies
$|A+B|\;\geq\;|A|+|B|-|H|,$
where
$H=\operatorname{Stab}(A+B)=\{\,g\in G:(A+B)+g=A+B\,\}$
is the stabilizer of $A+B$.
Let $G$ be an abelian group, $x,a\in G$ and let $k$ be a positive integer. A progression of length $k$ with difference $x$ and initial term $a$ is a subset of $G$ of the form $\{a,a+x,a+2x,...,a+(k-1)x\}$. We say $A$ is a semi-progression if $A\setminus\{a\}$ is a progression, for some $a\in A$. In the following lemma due to Kemperman, a certain family of critical pairs is characterized. Note that a pair $(A,B)$ of finite subsets of a group $G$ is called critical if |G|>|A+B|=|A|+|B|1|G|>|A+B|=|A|+|B|-1.
Lemma 2.6.
[16] Let $(A,B)$ be a critical pair of a finite abelian group $G$ with $|A|>1$, $|B|>1$ and $|A|+|B|-1\leq p(G)-2$. Then $A$ and $B$ are progressions with the same difference.
The following classical theorem was proved by Kemperman [15].
Theorem 2.7.
[15] Let $A$ and $B$ be nonempty finite subsets of a group $G$. Assume there exists an element $c\in A+B$ appearing exactly once as a sum $c=a+b$ with $a\in A,b\in B$. Then
$|A+B|\geq|A|+|B|-1.$
The following follows from Theorem 2.7.
Lemma 2.8.
Let $A$ and $B$ be nonempty finite subsets of a group $G$ such that $A$, $B$ and $A+B$ are all contained in a subset $X$ of $G\setminus\{0\}$. Then
$|X|\geq|A|+|B|+1.$
Proof.
Define $A_{0}=A\cup\{0\}$ and $B_{0}=B\cup\{0\}$. Then $0\in A_{0}+B_{0}$ and appears exactly once as a sum in $A_{0}+B_{0}$. That is, if $0=a+b$ with $a\in A$ and $b\in B$, since $0\notin A+B$, then $a=0$ or $b=0$, and hence $a=b=0$. Applying Theorem 2.7 to $A_{0}$ and $B_{0}$ we have
$|A_{0}+B_{0}|\geq|A_{0}|+|B_{0}|-1.$
Set $X=(A+B)\cup A\cup B$. Since $|A_{0}|=|A|+1$, $|B_{0}|=|B|+1$ and $A_{0}+B_{0}=(A+B)\cup A\cup B$, we have:
$|X|=|A_{0}+B_{0}|\geq|A|+|B|+1,$
as claimed. ∎
3. Main Results
Throughout, all matroids are assumed to be loopless. In this section, we investigate matchability within paving matroids. Although paving matroids possess a seemingly restrictive structure, Crapo and Rota [8] conjectured, and it is now widely believed, that asymptotically almost all matroids are paving. While this conjecture remains unresolved, it is supported by strong asymptotic estimates [21, 22]. What is certain, however, is that paving matroids constitute a large and significant class of matroids.
Note that matchability in the group setting can be viewed as a special case of matchability in the matroid setting. Let $A$ and $B$ be two finite nonempty subsets of an abelian group $G$, each of cardinality $n$, and assume $0\notin B$. Define a uniform matroid $M\cong U_{n,n}$ on the ground set $A$, and similarly define $N\cong U_{n,n}$ on the ground set $B$. Then $A$ is matched to $B$ in the group-theoretic sense if and only if $M$ is matched to $N$ in the matroid-theoretic sense. Note that uniform matroids form a very narrow subclass of all matroids, and $U_{n,n}$ represents a particularly special case.
The importance of paving matroids is not only that they extend sparse paving matroids, but that they make the structural source of matchability more visible. Since the classical group-theoretic notion of matching is recovered in the uniform case, paving matroid matchability extends the original problem of matchings within groups. Passing to all paving matroids, therefore, helps separate features of matchability that depend on near-uniformity from those that persist in a richer dependence structure. Our results show that, in paving matroids, matchability is governed by geometric features such as hyperplane intersections, hyperplane nullity, and stressed hyperplanes. This both explains why the theory extends beyond the sparse paving case and places the classical group case within a broader structural framework. Since paving matroids can also be related back to uniform matroids through relaxation (Theorem 3.23), the group-theoretic setting appears not as an isolated phenomenon, but as the first instance of a more general theory of matchings.
3.1. Matching Paving Matroids over Abelian Groups (Symmetric Case)
The following result was proved in [6] for matchings in sparse paving matroids under the symmetric case, providing a matroidal version of Theorem 1.1.
Theorem 3.1.
Let $M$ be a sparse paving matroid over $G$. Then $M$ is matched to itself if and only if $0\notin E(M)$.
We now extend this to the paving matroid analogue of Theorem 3.1.
Theorem 3.2.
Let $M$ be a paving matroid over $G$. Then $M$ is matched to itself if and only if $0\notin E(M)$.
Note that, for any pair of matroids $M,N$, the conditions $0\notin E(N)$ and $r(M)=r(N)$ are necessary for $M$ to be matched to $N$ (see [6, Proposition 1.7]). Thus, the “only if” implication in Theorem 3.2 is immediate. We now establish the reverse implication.
Proof.
We assume that $M$ is not a free matroid (the case when $M$ is free is addressed in Remark 3.5). Since $0\notin E(M)$, it follows from Theorem 1.1 that $E(M)$ is matched to itself in the group sense. That is, there exists a bijection $f:E(M)\to E(M)$ with $a+f(a)\notin E(M)$ for all $a\in E(M)$.
Let $\mathcal{M}=\{a_{1},\dots,a_{n}\}$ be a basis of $M$. Set $f(a_{i}):=b_{i}$ and define $\mathcal{N}=\{b_{1},\dots,b_{n}\}$. If $\mathcal{N}$ is a basis of $M$, then $\mathcal{M}$ is matched to $\mathcal{N}$ in the matroid sense and we are done. Otherwise, suppose that $\mathcal{N}$ is not a basis of $M$.
Claim 3.3.
There exist $x\in E(M)\setminus\mathcal{N}$ and $i\in[n]$ such that $a_{i}+x\notin E(M)$.
Proof of Claim.
Suppose not. Then
$\mathcal{M}+(E(M)\setminus\mathcal{N})\subseteq E(M).$
Since the three sets $\mathcal{M}$, $E(M)\setminus\mathcal{N}$, and $\mathcal{M}+(E(M)\setminus\mathcal{N})$ are contained in $E(M)$ and $0\notin E(M)$, Lemma 2.8 yields
$|E(M)|\;\geq\;|\mathcal{M}|+|E(M)\setminus\mathcal{N}|+1\;=\;|E(M)|+1,$
which is a contradiction. This completes the proof of the claim. ∎
Let $x$ and $i$ be as given by Claim 3.3. Without loss of generality, assume $i=n$ and define
$\mathcal{P}=\{b_{1},\dots,b_{n-1},x\}.$
Claim 3.4.
$\mathcal{P}$ is a basis of $M$.
Proof of the Claim.
Suppose not. Since $M$ is a paving matroid, we must then have $r(\mathcal{P})=n-1$. Consider the hyperplanes $H_{1}=\mathrm{cl}(\mathcal{N})$ and $H_{2}=\mathrm{cl}(\mathcal{P})$. By Theorem 2.1, it follows that
$|H_{1}\cap H_{2}|\leq n-2.$
However, it is clear that $|H_{1}\cap H_{2}|\geq n-1$, since $b_{i}\in H_{1}\cap H_{2}$ for all $i\in[n-1]$. This contradiction proves that $\mathcal{P}$ is indeed a basis. ∎
Finally, observe that the basis $\mathcal{M}$ is matched to the basis $\mathcal{P}$. Hence $M$ is matched to itself, completing the proof.
Remark 3.5.
In the proof of Theorem 3.2, if $M$ is a free matroid, then it has a unique basis, namely $E(M)$. Since $0\notin E(M)$, Theorem 1.1 guarantees that $E(M)$ is matched to itself in the group sense. Consequently, the basis $E(M)$ is also matched to itself in the matroid sense. Hence $M$ is matched to itself.
Remark 3.6.
The significance of Theorem 3.2 is that it shows the classical self-matching criterion from abelian groups (Theorem 1.1) is not confined to sets or to uniform matroids. Even in a paving matroid, where substantial dependence may occur, the only obstruction to symmetric self-matchability remains the trivial additive obstruction $0\in E(M)$. Thus, the theorem identifies the paving condition as a structural setting in which matroidal dependence does not create new symmetric matching obstructions. In particular, it shows that the source of self-matchability lies not merely in near-uniformity, but in the hyperplane geometry of paving matroids.
Example 3.7.
We construct a paving matroid over $\mathbb{Z}$ whose ground set does not contain 0, and then show that it is matched to itself in view of Theorem 3.2. To illustrate a case not covered by Theorem 3.1, we define a paving matroid that is not sparse paving.
Consider the matroid $M$ of rank $3$ with ground set $E(M)=[5]$ and bases
$\mathcal{B}(M)=\bigl\{\{i,j,5\}\mid i,j\in[4],\ i\neq j\bigr\}.$
The $3$-circuits are
$\mathcal{C}_{3}(M)\;=\;\big\{\{1,2,3\},\ \{1,2,4\},\ \{1,3,4\},\ \{2,3,4\}\big\}.$
Thus $M$ is a paving matroid of rank $3$. However, it is not sparse paving, since the subset $\{1,2,3\}$ is a $3$-subset that is neither a basis nor a hyperplane.
We use the construction of Theorem 3.2 to show that $M$ is matched to itself. First, consider the matching
[…]
in the group setting.
  • Consider the basis $\mathcal{M}=\{1,2,5\}$. Then
    $\mathcal{N}=f(\mathcal{M})=\{5,4,1\},$
    which is a basis of $M$, and $\mathcal{M}$ is matched to $\mathcal{N}$.
  • Consider the basis $\mathcal{M}=\{1,3,5\}$. Then
    $\mathcal{N}=f(\mathcal{M})=\{5,3,1\},$
    which is a basis of $M$, and $\mathcal{M}$ is matched to $\mathcal{N}$.
  • Consider the basis $\mathcal{M}=\{1,4,5\}$. Then
    $\mathcal{N}=f(\mathcal{M})=\{5,2,1\},$
    which is a basis of $M$, and $\mathcal{M}$ is matched to $\mathcal{N}$.
  • Consider the basis $\mathcal{M}=\{2,3,5\}$. Then
    $\mathcal{N}=f(\mathcal{M})=\{4,3,1\},$
    which is not a basis of $M$. Choose $a_{3}=5$ and $x=5$ as in Claim 3.3, and form the basis
    $\mathcal{P}=\{4,3,5\}$
    of $M$. Then $\mathcal{M}$ is matched to $\mathcal{P}$.
  • Consider the basis $\mathcal{M}=\{2,4,5\}$. Then
    $\mathcal{N}=f(\mathcal{M})=\{4,2,1\},$
    which is not a basis of $M$. Choose $a_{1}=2$ and $x=5$ as in Claim 3.3, and form the basis
    $\mathcal{P}=\{5,2,1\}$
    of $M$. Then $\mathcal{M}$ is matched to $\mathcal{P}$.
  • Consider the basis $\mathcal{M}=\{3,4,5\}$. Then
    $\mathcal{N}=f(\mathcal{M})=\{3,2,1\},$
    which is not a basis of $M$. Choose $a_{2}=4$ and $x=5$ as in Claim 3.3, and form the basis
    $\mathcal{P}=\{3,5,1\}$
    of $M$. Then $\mathcal{M}$ is matched to $\mathcal{P}$.
Since every basis $\mathcal{M}$ of $M$ is matched to a basis of $M$, it follows that $M$ is matched to itself.
Let $E$ be a nonempty set. A matroid basis system is a nonempty family $\mathcal{B}\subseteq 2^{E}$ of subsets, called bases, that satisfies the exchange axiom: for all distinct $B,B^{\prime}\in\mathcal{B}$ and every $x\in B\setminus B^{\prime}$, there exists $y\in B^{\prime}\setminus B$ such that $(B\setminus\{x\})\cup\{y\}\in\mathcal{B}$. It follows that $\mathcal{B}$ is the set of bases of a matroid (see Kung’s survey [17]).
Example 3.8.
In this example we give a non-paving matroid over an abelian group for which the construction used in the proof of Theorem 3.2 breaks down.
Let $G=\mathbb{Z}$ and let
$E=\{1,3,7,9\}.$
Define a matroid $M$ of rank $3$ on $E$ by declaring its bases to be exactly
$\mathcal{B}(M)=\bigl\{\{1,7,9\},\{3,7,9\}\bigr\}.$
It is immediate that this is a matroid: the collection $\mathcal{B}(M)$ is nonempty, and the basis exchange axiom is trivially satisfied since the only two bases differ by exchanging $1$ and $3$.
Since the $2$-subset $\{1,3\}$ is contained in no basis, it is dependent. Hence $M$ is not paving.
Now define a matching $f:E\to E$ in the group setting by
$f(1)=9,\qquad f(3)=1,\qquad f(7)=3,\qquad f(9)=7.$
Consider the basis
$\mathcal{M}=\{3,7,9\}.$
Following the construction in the proof of Theorem 3.2, set
$\mathcal{N}=f(\mathcal{M})=\{1,3,7\}.$
Then $\mathcal{N}$ is not a basis of $M$, since $\mathcal{N}\notin\mathcal{B}(M)$.
Now take $a_{3}=9$ and $x=9\in E\setminus\mathcal{N}$. Since
$9+9=18\notin E,$
the analogue of Claim 3.3 holds. Replacing the corresponding element $7$ of $\mathcal{N}$ by $x=9$, we obtain
$\mathcal{P}=\{1,3,9\}.$
Again, $\mathcal{P}$ is not a basis of $M$, since $\mathcal{P}\notin\mathcal{B}(M)$.
Therefore, the replacement step used in the proof of Theorem 3.2 may fail for non-paving matroids. This shows that the paving assumption is needed in order to conclude that the modified set is again a basis.
3.2. Matching Paving Matroids over Abelian Groups (Asymmetric Case)
The following result about the matchability of sparse paving matroids in the asymmetric situation was proved in [6].
Theorem 3.9.
Let $M$ be a matroid over $G$ and let $N$ be a sparse paving matroid over $G$, both of the same rank $n$, with $0\notin E(N)$. Assume further that one of the following conditions holds:
  1. (1)
    $|E(M)|<\min\{|E(N)|-1,\,p(G)\}$, or
  2. (2)
    $E(M)$ is not a progression, $G$ is finite, |E(M)|=|E(N)|1|E(M)|=|E(N)|-1, and $|E(N)|<p(G)$, or
  3. (3)
    $E(M)$ is neither a progression nor a semi-progression, $G$ is finite, $|E(M)|=|E(N)|<p(G)$, or
  4. (4)
    |E(M)|<|E(N)|n1|E(M)|<|E(N)|-n-1.
Then $M$ is matched to $N$.
We now extend this result to general paving matroids by incorporating the hyperplane-nullity parameter. In particular, we generalize the sparse paving case ($t\leq 1$) to arbitrary paving matroids, with explicit dependence on the parameter $t$.
Theorem 3.10.
Let $M$ be a matroid over $G$ and let $N$ be a paving matroid over $G$, both of the same rank $n$, with $0\notin E(N)$. Let $t=\mathrm{null}(\mathcal{H}_{N})$ with $0<t\leq n$. Assume further that one of the following conditions holds:
  1. (1)
    $|E(M)|<\min\{|E(N)|-2t+1,\,p(G)\}$, or
  2. (2)
    $E(M)$ is not a progression, $G$ is finite, |E(M)|=|E(N)|2t+1|E(M)|=|E(N)|-2t+1, and $|E(N)|<p(G)$, or
  3. (3)
    $E(M)$ is neither a progression nor a semi-progression, $G$ is finite, |E(M)|=|E(N)|2t+2|E(M)|=|E(N)|-2t+2, and $|E(N)|<p(G)$, or
  4. (4)
    |E(M)|<|E(N)|n1|E(M)|<|E(N)|-n-1.
Then $M$ is matched to $N$.
Proof of Theorem 3.10 (1)-(3).
We begin by addressing the following three special cases:
  • $M$ is a free matroid: In this case, $\mathcal{M}=E(M)$ is the unique basis of $M$. Let $\mathcal{N}$ be a basis of $N$. By Theorem 1.3, $\mathcal{M}$ is matched to $\mathcal{N}$ in the group-theoretic sense; that is, there exists a bijection $f:\mathcal{M}\to\mathcal{N}$ such that $a+f(a)\notin\mathcal{M}$ for all $a\in\mathcal{M}$. This implies that $\mathcal{M}$ is also matched to $\mathcal{N}$ in the matroid sense. Hence, $M$ is matched to $N$.
  • $t=1$: By Lemma 2.3, NN is a sparse paving matroid. Then, by Theorem 3.9, $M$ is matched to $N$.
  • $n=1$: Let $\mathcal{M}=\{a\}$ be a basis of $M$. Since $|E(M)|<|E(N)|$, there exists an element $b\in E(N)$ such that $a+b\notin E(M)$. Define $\mathcal{N}=\{b\}$. As $N$ is loopless, $\mathcal{N}$ is a basis of $N$. Thus, $\mathcal{M}$ is matched to $\mathcal{N}$, and therefore $M$ is matched to $N$.
We now assume that $M$ is not a free matroid, that $t>1$, and that $n>1$. Suppose that one of the conditions (1)-(3) holds. Then $|E(M)|<p(G)$. Choose a subset $A\subseteq E(N)$ such that $|A|=|E(M)|$. Since $|A|=|E(M)|<p(G)$, Theorem 1.3 guarantees the existence of a matching $f:E(M)\to A$ in the group-theoretic sense.
Let $\mathcal{M}=\{a_{1},\ldots,a_{n}\}$ be a basis of $M$, and define $b_{i}:=f(a_{i})$ for all $i\in[n]$. Set $\mathcal{N}=\{b_{1},\ldots,b_{n}\}\subseteq E(N)$. Since $N$ is a paving matroid of rank $n$, the set $\mathcal{N}$ is either a basis or a circuit.
If $\mathcal{N}$ is a basis, then $\mathcal{M}$ is matched to $\mathcal{N}$ in the matroid sense, and hence $M$ is matched to $N$, completing the proof in this case. If instead $\mathcal{N}$ is a circuit, we proceed by proving the following claim.
Claim 3.11.
For every integer $s$ with $s\in[t]$, there exist distinct elements $x_{1},\dots,x_{s}\in E(N)\setminus\mathcal{N}$ and distinct indices $i_{1},\dots,i_{s}\in[n]$ such that
$a_{i_{j}}+x_{j}\notin E(M)\qquad\text{for every }j\in[s].$
In particular, the conclusion holds for $s=t$.
Proof of the Claim.
We proceed by induction on $s$.
Base case ($s=1$). We show that there exist $x\in E(N)\setminus\mathcal{N}$ and $a\in\mathcal{M}$ such that $a+x\notin E(M)$. Suppose, for contradiction, that no such pair $(a,x)$ exists. Then
$\mathcal{M}+(E(N)\setminus\mathcal{N})\subseteq E(M).$
Applying Theorem 2.5 to the sets $\mathcal{M}$ and $E(N)\setminus\mathcal{N}$, we obtain a subgroup $H\leq G$ such that
(2) $|\mathcal{M}+(E(N)\setminus\mathcal{N})|$ $\geq|\mathcal{M}|+|E(N)\setminus\mathcal{N}|-|H|,$
(3) $\mathcal{M}+(E(N)\setminus\mathcal{N})+H$ $=\mathcal{M}+(E(N)\setminus\mathcal{N}).$
Since $\mathcal{M}+(E(N)\setminus\mathcal{N})\subseteq E(M)$, it follows from (3) that
$|H|\leq|\mathcal{M}+(E(N)\setminus\mathcal{N})+H|=|\mathcal{M}+(E(N)\setminus\mathcal{N})|\leq|E(M)|<p(G),$
so $H=\{0\}$. Substituting this into (2) yields
(4) $|E(M)|\;\geq\;|\mathcal{M}|+|E(N)\setminus\mathcal{N}|-1=n+(|E(N)|-n)-1=|E(N)|-1>|E(N)|-2t+1.$
We now consider three separate cases corresponding to conditions (1), (2), and (3) of the theorem.
Case (1). The condition
$|E(M)|<\min\{|E(N)|-2t+1,\,p(G)\}$
directly contradicts inequality (4).
Case (2). The assumption |E(M)|=|E(N)|2t+1|E(M)|=|E(N)|-2t+1 directly contradicts inequality (4).
Case (3). Assume that $G$ is finite. Combining the assumption |E(M)|=|E(N)|2t+2|E(M)|=|E(N)|-2t+2 with inequality (4), we obtain
$|E(M)|\;\geq\;|\mathcal{M}+(E(N)\setminus\mathcal{N})|\;\geq\;|\mathcal{M}|+|E(N)\setminus\mathcal{N}|-1\;\geq\;|E(N)|-2t+2=|E(M)|.$
Therefore,
$|\mathcal{M}|+|E(N)\setminus\mathcal{N}|-1\;=\;|\mathcal{M}+(E(N)\setminus\mathcal{N})|.$
The pair $(\mathcal{M},\,E(N)\setminus\mathcal{N})$ is critical and satisfies the conditions of Theorem 2.6 because:
  • Since $n>1$, we have $|\mathcal{M}|>1$.
  • $|E(N)\setminus\mathcal{N}|=|E(N)|-n>1$.
  • $|\mathcal{M}|+|E(N)\setminus\mathcal{N}|-1=|E(N)|-1\leq p(G)-2$.
Hence, both sets are progressions with the same common difference. Since
$E(M)=\mathcal{M}+(E(N)\setminus\mathcal{N}),$
it follows that $E(M)$ is also a progression, a contradiction.
Inductive hypothesis. Assume that distinct elements $x_{1},\dots,x_{s-1}\in E(N)\setminus\mathcal{N}$ and distinct indices $i_{1},\dots,i_{s-1}\in[n]$ have been chosen such that
$a_{i_{j}}+x_{j}\notin E(M)\quad\text{for every }j\in[s-1].$
Inductive step. We show that there exists
$x_{s}\in E(N)\setminus(\mathcal{N}\cup\{x_{1},\dots,x_{s-1}\})\quad\text{and}\quad i_{s}\in[n]\setminus\{i_{1},\dots,i_{s-1}\}$
such that
$a_{i_{s}}+x_{s}\notin E(M).$
Suppose, for contradiction, that no such pair exists. Then
(5) $(E(N)\setminus(\mathcal{N}\cup\{x_{1},\dots,x_{s-1}\}))+(\mathcal{M}\setminus\{a_{i_{1}},\dots,a_{i_{s-1}}\})\subseteq E(M).$
Applying Theorem 2.5 to the sets $\mathcal{M}\setminus\{a_{i_{1}},\dots,a_{i_{s-1}}\}$ and $E(N)\setminus(\mathcal{N}\cup\{x_{1},\dots,x_{s-1}\})$, we obtain a subgroup $H\leq G$ such that
(6) $\Big|(\mathcal{M}\setminus\{a_{i_{1}},\ldots,a_{i_{s-1}}\})+(E(N)\setminus(\mathcal{N}\cup\{x_{1},\ldots,x_{s-1}\}))\Big|$
$\qquad\geq|\mathcal{M}\setminus\{a_{i_{1}},\ldots,a_{i_{s-1}}\}|+|E(N)\setminus(\mathcal{N}\cup\{x_{1},\ldots,x_{s-1}\})|-|H|,$
$(\mathcal{M}\setminus\{a_{i_{1}},\ldots,a_{i_{s-1}}\})+(E(N)\setminus(\mathcal{N}\cup\{x_{1},\ldots,x_{s-1}\}))+H$
$\qquad=(\mathcal{M}\setminus\{a_{i_{1}},\ldots,a_{i_{s-1}}\})+(E(N)\setminus(\mathcal{N}\cup\{x_{1},\ldots,x_{s-1}\})).$
In a similar manner as in the base case, we may argue that $H=\{0\}$. Combining this with inequalities (5) and (6), we conclude that
(7) $|E(M)|$ $\geq\big|(\mathcal{M}\setminus\{a_{i_{1}},\ldots,a_{i_{s-1}}\})+(E(N)\setminus(\mathcal{N}\cup\{x_{1},\ldots,x_{s-1}\}))\big|$
$\geq|\mathcal{M}\setminus\{a_{i_{1}},\ldots,a_{i_{s-1}}\}|+|E(N)\setminus(\mathcal{N}\cup\{x_{1},\ldots,x_{s-1}\})|-1$
$=|E(N)|-2s+1\geq|E(N)|-2t+1.$
We now consider three separate cases corresponding to conditions (1), (2), and (3) of the theorem.
Case (1). The condition
$|E(M)|<\min\{|E(N)|-2t+1,\,p(G)\}$
contradicts inequality (7).
Case (2). Assume |E(M)|=|E(N)|2t+1|E(M)|=|E(N)|-2t+1. From (7) we conclude that
$\Big|(\mathcal{M}\setminus\{a_{i_{1}},\ldots,a_{i_{s-1}}\})+\big(E(N)\setminus(\mathcal{N}\cup\{x_{1},\ldots,x_{s-1}\})\big)\Big|\\ =\big|\mathcal{M}\setminus\{a_{i_{1}},\ldots,a_{i_{s-1}}\}\big|+\big|E(N)\setminus(\mathcal{N}\cup\{x_{1},\ldots,x_{s-1}\})\big|-1.$
Therefore
$\big(\mathcal{M}\setminus\{a_{i_{1}},\ldots,a_{i_{s-1}}\},\,E(N)\setminus(\mathcal{N}\cup\{x_{1},\ldots,x_{s-1}\})\big)$
is a critical pair satisfying the conditions of Theorem 2.6, since:
  • $|\mathcal{M}\setminus\{a_{i_{1}},\ldots,a_{i_{s-1}}\}|=n-(s-1)\geq n-(t-1)>1$;
  • As $M$ is not a free matroid and $t>1$, we have
    $|E(N)\setminus(\mathcal{N}\cup\{x_{1},\ldots,x_{s-1}\})|$ $=|E(N)|-(n+s-1)$
    $\geq|E(N)|-(n+t-1)$
    $=|E(M)|+2t-1-(n+t-1)$
    $>t>1.$
  • $|\mathcal{M}\setminus\{a_{i_{1}},\ldots,a_{i_{s-1}}\}|$ $+|E(N)\setminus(\mathcal{N}\cup\{x_{1},\ldots,x_{s-1}\})|-1$
    $=|E(M)|$
    $\leq|E(N)|-1$
    $\leq p(G)-2.$
By Theorem 2.6, both sets are progressions with the same common difference. Thus
$E(M)=(\mathcal{M}\setminus\{a_{i_{1}},\ldots,a_{i_{s-1}}\})+\big(E(N)\setminus(\mathcal{N}\cup\{x_{1},\ldots,x_{s-1}\})\big)$
is itself a progression, contradicting the assumption that $E(M)$ is not a progression.
Case (3). Assume |E(M)|=|E(N)|2t+2|E(M)|=|E(N)|-2t+2. From (7) it follows that one of the following must hold:
$|\mathcal{M}\setminus\{a_{i_{1}},\ldots,a_{i_{s-1}}\}|+|E(N)\setminus(\mathcal{N}\cup\{x_{1},\ldots,x_{s-1}\})|-1=|E(M)|,$
or
$|\mathcal{M}\setminus\{a_{i_{1}},\ldots,a_{i_{s-1}}\}|+|E(N)\setminus(\mathcal{N}\cup\{x_{1},\ldots,x_{s-1}\})|-1=|E(M)|-1.$
We now split into two subcases:
  1. (1)
    If
    $\big|\mathcal{M}\setminus\{a_{i_{1}},\ldots,a_{i_{s-1}}\}\big|+\big|E(N)\setminus(\mathcal{N}\cup\{x_{1},\ldots,x_{s-1}\})\big|-1=|E(M)|,$
    then, as in Case (2), the pair
    $\big(\mathcal{M}\setminus\{a_{i_{1}},\ldots,a_{i_{s-1}}\},\,E(N)\setminus(\mathcal{N}\cup\{x_{1},\ldots,x_{s-1}\})\big)$
    is critical and satisfies the conditions of Theorem 2.6. Hence both $\mathcal{M}\setminus\{a_{i_{1}},\ldots,a_{i_{s-1}}\}$ and $E(N)\setminus(\mathcal{N}\cup\{x_{1},\ldots,x_{s-1}\})$ are progressions with the same common difference. Since
    $E(M)=(\mathcal{M}\setminus\{a_{i_{1}},\ldots,a_{i_{s-1}}\})+(E(N)\setminus(\mathcal{N}\cup\{x_{1},\ldots,x_{s-1}\})),$
    it follows that $E(M)$ is itself a progression, a contradiction.
  2. (2)
    If
    $\big|\mathcal{M}\setminus\{a_{i_{1}},\ldots,a_{i_{s-1}}\}\big|+\big|E(N)\setminus(\mathcal{N}\cup\{x_{1},\ldots,x_{s-1}\})\big|-1=|E(M)|-1,$
    then $|E(M)|=|E(N)|$. On the other hand, we also know that |E(M)|=|E(N)|2t+2|E(M)|=|E(N)|-2t+2, which forces $t=1$, a contradiction.
Thus the proof of Claim 3.11 is complete. ∎
Now let $x_{1},\dots,x_{t}$ and $i_{1},\dots,i_{t}$ be obtained as in Claim 3.11. Without loss of generality, we may assume that $i_{j}=j$ for every $j\in[t]$. That is, $a_{j}+x_{j}\notin E(M)$ for all $j\in[t]$. Set $X=\{x_{1},\dots,x_{t}\}$ and $\mathcal{P}=X\cup\mathcal{N}$.
Claim 3.12.
$r(\mathcal{P})=n$.
Proof of the Claim.
Assume for contradiction that $r(\mathcal{P})\neq n$. Since $N$ is a paving matroid and $|\mathcal{P}|>n$, it follows that $r(\mathcal{P})=n-1$. Let $H=\mathrm{cl}(\mathcal{P})$. Then $H$ is a hyperplane, and we have
$t=\mathrm{null}(\mathcal{H}_{N})\geq\mathrm{null}(H)=|H|-(n-1)\geq|\mathcal{P}|-(n-1)=(n+t)-(n-1)=t+1,$
a contradiction. Hence $r(\mathcal{P})=n$. ∎
We now define $\mathcal{N}_{0}=\mathcal{N}$ and consider the following $n$-subsets of $E(N)$:
$\mathcal{N}_{i}=\{x_{1},\dots,x_{i},b_{i+1},\dots,b_{n}\},\quad\text{for all }i\in[t].$
Claim 3.13.
$\mathcal{N}_{i}$ is a basis of $N$ for some $i\in[t]$.
Proof of the Claim.
Suppose, for contradiction, that every $\mathcal{N}_{i}$ is not a basis. Since $N$ is a paving matroid, this means $r(\mathcal{N}_{i})=n-1$. Set $H_{i}=\mathrm{cl}(\mathcal{N}_{i})$ for $0\leq i\leq t$. Clearly, each $H_{i}$ is a hyperplane of $N$. On the other hand,
$|H_{i}\cap H_{i+1}|\;\geq\;|\mathcal{N}_{i}\cap\mathcal{N}_{i+1}|=n-1.$
Invoking Theorem 2.1, we conclude that $H_{i}=H_{i+1}$ for all $0\leq i\leq t-1$. Thus,
$\mathcal{P}\subseteq\bigcup_{i=0}^{t}H_{i}=H_{0},$
implying $r(\mathcal{P})\leq r(H_{0})=n-1$. This contradicts the fact that $r(\mathcal{P})=n$. Hence, $\mathcal{N}_{i}$ is a basis of $N$ for some $i\in[t]$.
Assume that $i\in[t]$ is obtained as in Claim 3.13. Then $\mathcal{M}$ is matched to $\mathcal{N}_{i}$ via the assignment
$a_{j}\mapsto x_{j}\quad\text{for }1\leq j\leq i,\qquad a_{k}\mapsto b_{k}\quad\text{for }i+1\leq k\leq n.$
Therefore $M$ is matched to $N$. ∎
Remark 3.14.
It is worth noting that, although the conditions in Theorem 3.10-(3), such as $E(M)$ not being a semi-progression may seem unused in the proof of Theorem 3.10, they are in fact applied in the special case $t=1$. In this instance, we invoke Theorem 3.9, where this condition plays a role.
Next, we complete the proof of Theorem 3.10-(4). In this case, the condition $|E(N)|<p(G)$ is not imposed. Hence, unlike parts (1)–(3), where Theorem 1.3 is applicable, we cannot invoke it here. To establish the existence of matchings in the group sense (and subsequently in the matroid sense), we instead employ elementary techniques from additive number theory, followed by standard matroidal arguments.
Proof of Theorem 3.10 (4).
Let $\mathcal{M}=\{a_{1},\cdots,a_{n}\}$ be a basis for $M$. We show that $\mathcal{M}$ is matched to a basis of $N$.
Claim 3.15.
There exist $n$ distinct elements $b_{1},b_{2},\dots,b_{n}\in E(N)$ such that
$a_{i}+b_{i}\notin E(M)\quad\text{for all }i\in[n].$
Proof of the Claim.
We first choose $b_{1}\in E(N)$ so that $a_{1}+b_{1}\notin E(M)$. Such $b_{1}$ exists since $|E(M)|<|E(N)|$. Next, choose $b_{2}\in E(N)\setminus\{b_{1}\}$ such that $a_{2}+b_{2}\notin E(M)$. Such $b_{2}$ exists as |E(M)|<|E(N)|1|E(M)|<|E(N)|-1.
We proceed by induction on $n$. Suppose $b_{1},\dots,b_{j}$ have been chosen. We show that there exists $b_{j+1}\in E(N)\setminus\{b_{1},\dots,b_{j}\}$ satisfying $a_{j+1}+b_{j+1}\notin E(M)$. Assume to the contrary that no such $b_{j+1}$ exists. Then
$a_{j+1}+\big(E(N)\setminus\{b_{1},\dots,b_{j}\}\big)\subseteq E(M),$
which implies
$|E(N)\setminus\{b_{1},\dots,b_{j}\}|\leq|E(M)|.$
Thus
$|E(M)|\geq|E(N)|-j\geq|E(N)|-n,$
a contradiction. This proves the claim. ∎
Let $\mathcal{N}=\{b_{1},\dots,b_{n}\}$, where $b_{1},\dots,b_{n}$ are as above. Since $N$ is a paving matroid, then $\mathcal{N}$ is either a basis of $N$ or a circuit of rank n1n-1. If $\mathcal{N}$ is a basis, then $\mathcal{M}$ is matched to $\mathcal{N}$ and we are done. Suppose $\mathcal{N}$ is a circuit of rank n1n-1.
Claim 3.16.
There exist $x\in E(N)\setminus\mathcal{N}$ and $i\in[n]$ such that $a_{i}+x\notin E(M)$.
Proof of Claim.
Assume to the contrary that no such $x$ and $a_{i}$ exist. Then
$M+\big(E(N)\setminus\{b_{1},\dots,b_{n}\}\big)\subseteq E(M),$
which implies
$|E(N)|-n=\big|E(N)\setminus\{b_{1},\dots,b_{n}\}\big|\leq|E(M)|.$
This contradicts the assumption |E(M)|<|E(N)|n1|E(M)|<|E(N)|-n-1. ∎
Choose $i\in[n]$ and $x\in E(N)\setminus\mathcal{N}$ as in Claim 3.16. Consider the subset
$\mathcal{P}=\{b_{1},\dots,b_{i-1},x,b_{i+1},\dots,b_{n}\}\subseteq E(N).$
In a similar manner as in the proof of Theorem 3.10 (parts (1)-(3)), one may argue that $\mathcal{P}$ is a basis of $N$. Evidently, $\mathcal{M}$ is matched to $\mathcal{P}$. Hence $M$ is matched to $N$, completing the proof.
Remark 3.17.
Note that Theorem 3.10 does not address the case $t=0$, since this forces NN to be a uniform matroid (see Corollary 2.4). Matchability in uniform matroids is straightforward and has been completely characterized in [6, Proposition 2.9].
Remark 3.18.
Note that in the asymmetric problem of the sparse paving case, dependence is controlled by circuit-hyperplanes, whereas in the general paving case, the size of the obstruction is measured by hyperplane nullity. Theorem 3.10 shows that the existence of matchings is governed not merely by the size gap |E(N)||E(M)||E(N)|-|E(M)|, but also by the extent to which $N$ deviates from uniformity, as quantified by the hyperplane nullity parameter $t$.
Example 3.19.
We define two matroids $M$ and $N$ over the cyclic group $\mathbb{Z}_{13}$.
Let $E(M)=\{\overline{2},\overline{4},\overline{6},\overline{11}\}$ and define matroid $M$ of rank 3 by its set of bases
$\mathcal{B}(M)=\binom{E(M)}{3}\setminus\big\{\{\overline{2},\overline{4},\overline{6}\}\big\}.$
Equivalently, the unique $3$-circuit is $\{\overline{2},\overline{4},\overline{6}\}$, and the bases are $\{\overline{2},\overline{4},\overline{11}\}$, $\{\overline{2},\overline{6},\overline{11}\}$, and $\{\overline{4},\overline{6},\overline{11}\}$.
Let $E(N)=\{\overline{1},\ldots,\overline{9}\}$ and fix $H=\{\overline{1},\overline{2},\overline{3},\overline{4}\}\subseteq E(N)$. Define a rank-$3$ matroid $N$ by
$\mathcal{B}(N)=\binom{E(N)}{3}\setminus\binom{H}{3}$,
i.e., the only $3$-subsets that are not bases are the four triples inside $H$. Then $N$ is a paving matroid that is not sparse paving. Its largest hyperplane is $H$ (size $4$), so its hyperplane nullity is |H|r(H)=42=2|H|-r(H)=4-2=2. Thus $M$ and $N$ satisfy Condition (1) of Theorem 3.10 and $M$ must be matched to $N$.
Exhibiting matching bases. We now explicitly show that $M$ is matched to $N$ by pairing each basis of $M$ with a basis of $N$:
  • $\{\overline{2},\overline{4},\overline{11}\}$ in $M$ matches $\{\overline{1},\overline{3},\overline{7}\}$ in $N$;
  • $\{\overline{2},\overline{6},\overline{11}\}$ in $M$ matches $\{\overline{1},\overline{4},\overline{9}\}$ in $N$;
  • $\{\overline{4},\overline{6},\overline{11}\}$ in $M$ matches $\{\overline{1},\overline{6},\overline{9}\}$ in $N$.
3.3. Matchable Bases through Stressed Hyperplanes
Let $E$ be a nonempty set. Recall that A matroid basis system is a nonempty family $\mathcal{B}\subseteq 2^{E}$ of subsets, called bases, that satisfies the exchange axiom: for all distinct $B,B^{\prime}\in\mathcal{B}$ and every $x\in B\setminus B^{\prime}$, there exists $y\in B^{\prime}\setminus B$ such that $(B\setminus\{x\})\cup\{y\}\in\mathcal{B}$. It follows that $\mathcal{B}$ is the set of bases of a matroid.
Following [12], a hyperplane $H$ of a matroid $M$ is said to be stressed if every (r(M)1)(r(M)-1)-subset of $H$ is independent.
Proposition 3.20 ([12]).
A matroid is paving if and only if all its hyperplanes are stressed.
Definition 3.21.
Let $M$ be a matroid of rank $n=r(M)$ with bases $\mathcal{B}$, and let $S\subseteq E(M)$. Denote by $\binom{S}{n}$ the family of $n$-subsets of $S$. Suppose that no element of $\binom{S}{n}$ is a basis of $M$. We say that $S$ can be relaxed if
$\mathrm{Rel}_{S}(\mathcal{B})\;:=\;\mathcal{B}\,\cup\,\binom{S}{n}$
is again a matroid basis system. In this case, the resulting matroid is called the relaxation of $M$ at $S$, and is denoted by $\mathrm{Rel}_{S}(M)$.
Proposition 3.22 ([12]).
If $H$ is a stressed hyperplane of $M$, then $H$ can be relaxed.
As a consequence, given an $n$-rank paving matroid $M$ on $m$ elements, one can obtain the uniform matroid $U_{n,m}$ by repeatedly relaxing its stressed hyperplanes. Indeed, [12] shows that every relaxation of a paving matroid is again paving, and that relaxing all stressed hyperplanes of $M$ yields $U_{n,m}$.
Theorem 3.23.
Let $M$ be a matroid and $N$ a paving matroid over an abelian group $G$, both of rank $n$, with $0\notin E(N)$ and
$|E(M)|\;\leq\;\min\{\,p(G)-1,\,|E(N)|\,\}.$
If $\mathcal{M}$ is a basis of $M$, then $\mathcal{M}$ is either matched to a basis of $N$, or to a basis of $\mathrm{Rel}_{S}(N)$ for some $S\subseteq E(N)$.
Proof.
Let $A\subseteq E(N)$ with $|A|=|E(M)|$. By Theorem 1.3, $E(M)$ is matched to $A$ in the group sense; equivalently, there exists a matching $f:E(M)\to A$. Let $\mathcal{M}=\{a_{1},\dots,a_{n}\}$ and set $b_{i}=f(a_{i})$ for $i\in[n]$, and $\mathcal{N}=\{b_{1},\dots,b_{n}\}$.
If $\mathcal{N}$ is a basis of $N$, then $\mathcal{M}$ is matched to $\mathcal{N}$ in the matroid sense and we are done. Otherwise, set $H=\operatorname{cl}(\mathcal{N})$. Since $N$ is paving, $H$ is a hyperplane of $N$, and by Proposition 3.20 it is stressed. Hence, by Proposition 3.22, $H$ can be relaxed. By definition of relaxation, $\mathcal{N}$ is a basis of $\mathrm{Rel}_{H}(N)$ (equivalently, $\mathcal{N}\in\mathrm{Rel}_{H}(\mathcal{B}_{N})$, where $\mathcal{B}_{N}$ is the basis system of $N$). Consequently, $\mathcal{M}$ is matched to $\mathcal{N}$ in the matroid sense in $\mathrm{Rel}_{H}(N)$, as required. ∎
Remark 3.24.
When matchings between matroid bases are not available, stressed hyperplanes show that compatibility can often be recovered by relaxing independence constraints. A stressed hyperplane may be viewed as a minimal obstruction to uniformity. Relaxing such hyperplanes, turning certain dependent sets into bases, gradually transforms a paving matroid toward a uniform matroid while preserving structure. Theorem 3.23 shows that this process offers an alternative route to matchability: even if bases of $M$ cannot be matched to those of $N$ directly, they may be matched to bases of a relaxation $\mathrm{Rel}_{S}(N)$. Thus matchability is not a fragile property tied to exact independence, but a stable phenomenon that persists under natural modifications. Relaxation therefore links the restrictive setting of paving matroids with the uniform case, where matching problems are straightforward.
3.4. Concluding Remarks and Future Directions
  1. (1)
    Let $M$ be a matroid of rank $n$ and let $k\in[n]$. We say that $M$ is $k$-paving if every circuit of $M$ has size >nk>n-k. For $k=1$ this recovers the usual paving condition, where all circuits have size at least $n$. Clearly, the following hierarchy holds:
    $\text{paving}=1\text{-paving}\ \subseteq\ 2\text{-paving}\ \subseteq\ \cdots\ \subseteq\ n\text{-paving}=\{\text{all matroids of rank }n\}.$
    The notion of $k$-paving matroids was first introduced in [23], where a characterization was given; see also [24] for more recent developments. We believe that some of the techniques developed here for the matchability of paving matroids could be adapted to investigate matchability in the broader class of $k$-paving matroids for $k>1$.
    In particular, one may ask whether the dependence on $t$ in Theorem 3.10 extends in a clean way:
    $|E(M)|<|E(N)|-2t+1\;\;\longrightarrow\;\;|E(M)|<|E(N)|-c(k,t)+1,$
    for some explicit function $c(k,t)$. Further work in this direction appears promising.
  2. (2)
    For a paving matroid $N$, one may obtain the uniform matroid $U_{n,m}$ by successively relaxing its stressed hyperplanes, forming a hierarchy of relaxations
    $N\;\longrightarrow\;\mathrm{Rel}_{H_{1}}(N)\;\longrightarrow\;\;\cdots\;\longrightarrow\;U_{n,m}.$
    Assuming the conditions of Theorem 3.23, what is the minimal number of steps in this hierarchy needed to guarantee that $M$ can be matched to some consecutive relaxations of $N$? Clearly, this process must terminate, since in the worst-case scenario $M$ is eventually matched to $U_{n,m}$. Is a single relaxation always sufficient, or can certain pairs $(M,N)$ require multiple relaxations within the hierarchy?
Data sharing. Data sharing is not applicable to this article as no datasets were generated or analyzed.
Conflict of interest. To the best of our knowledge, no conflict of interest, whether of a financial or personal nature, has influenced the work presented in this article.